OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2

OpenStudy (anonymous):

Has no solution because they are not intersect... First, y = 1 - 2x Second , 4x + 2y = -2 divide by 2 2x + y = -1 => y = -2x -1 Look at tangents... Tangents are same...

OpenStudy (anonymous):

that's a choice, thanks!